# Fast and Robust Visuomotor Riemannian Flow Matching Policy

### v2

@@ V-A1 RFMP and SRFMP Implementation
Our RFMP implementation builds on the RFM framework from Chen and Lipman [16]. We parameterize the vector field $v_{t}(\bm{a}|\bm{o};\bm{\theta})$ using the UNet architecture employed in DP [2], which consists of $3$ layers with downsampling dimensions of $(256,512,1024)$ and $(128,256,512)$ for simulated and real-world tasks. Each layer employs a $1$-dimensional convolutional residual network as proposed in [29]. We implement a Feature-wise Linear Modulation (FiLM) [59] to incorporate the observation condition vector ${\bm{o}}$ and time step $t$ into the UNet. Instead of directly feeding the FM time step $t$ as a conditional variable, we first project it into a higher-dimensional space using a sinusoidal embedding module, similarly to DP. For tasks with image-based observations, we leverage the same vision perception backbone as in DP [2]. Namely, we use a standard ResNet-$18$ in which we replace: (1) the global average pooling with a spatial softmax pooling, and (2) BatchNorm with GroupNorm. Our SRFMP implementation builds on the SFM framework [20]. We implement the same UNet as RFMP to represent $v_{{\bm{x}}}$ by replacing the time step $t$ by the temperature parameter $\tau$. We introduce an additional Multi-Layer Perceptron (MLP) to learn $v_{\tau}$. As for $t$ in RFMP, we employed a sinusoidal embedding for the input $\tau$. The boundaries $\tau_{0}$ and $\tau_{1}$ are set to 00 and $1$ in all experiments.
@@
We implement different prior distributions for different tasks. For the Euclidean Push-T and the Robomimic tasks, the action space is $\mathbb{R}^{d}$ and we thus define the prior distribution as a Euclidean Gaussian distribution for both RFMP and SRFMP. For the Sphere Push-T, the action space is the hypersphere $\mathcal{S}^{2}$. In this case, we test two types of Riemannian prior distribution, namely a spherical uniform distribution, and a wrapped Gaussian distribution [60, 61], illustrated in Figure 5. For the Franka kitchen benchmark, we define the action space as the product of manifolds $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{2}$, whose components represent the end-effector position, the end-effector orientation (encoded as quaternions), and the gripper fingers position. Regarding the real robot tasks, the action space is defined similarly as the product of manifolds $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{1}$, whose components represent the position, orientation (encoded as quaternions), and opening of the gripper. For the Franka kitchen benchmark and the real robot tasks, the Euclidean and hypersphere parts employ Euclidean Gaussian distributions and a wrapped Gaussian distribution, respectively. Notice that this choice of prior distributions follows common practice in flow matching literature, where Gaussian and uniform priors are widely adopted due to their simplicity and effectiveness [16]. While more sophisticated approaches could be explored, for example using Gaussian Process as prior distributions [62], or learning task-dependent priors as in D-Flow [63], the choice depends on the problem at hand. We focus here on standard priors to maintain simplicity and consistency across tasks, while isolating their impact on our framework.
TABLE I: Hyperparameters for all experiments: Resolution of original and cropped image, number of parameters for the learned vector field using RFMP and SRFMP, number of ResNet parameters, training epochs and batch size.
Experiment Image res. Crop res. RFMP VF Num. params SRFMP VF Num. params ResNet params Epochs Batch size
Push-T Tasks
[…]
State-based simulation tasks
[…]
Vision-based simulation tasks
[…]
Real-world experiments
[…]
@@
Figure 5: 2D visualization of prior distributions used in the Sphere Push-T task: wrapped Gaussian distribution (left) and sphere uniform distribution (right). The sphere Gaussian distribution is obtained by first sampling from a Euclidean Gaussian distribution on a tangent space of the sphere (the red points), followed by the exponential map, which projects the samples onto the sphere manifold $\mathcal{S}^{d}$ (blue points). The spherical uniform distribution is computed by normalizing samples from a zero-mean Euclidean Gaussian distribution.
@@ V-A2 Baselines
TABLE II: Action space dimensionality for all experiments.
@@
Euclidean Push-T $16$ $2$ $32$
Sphere Push-T $16$ $2$ $32$
Robomimic tasks $16$ $7$ $112$
Franka Kitchen $216$ $9$ $1944$
Real-world experiments $16$ $7$ $112$
@@ V-A3 Evaluation methodology
We evaluate all policies using three key metrics: (1) The performance, computed as the average task-depending score across all trials, with $50$ trials for each simulated task, and $10$ trials for each real-world task; (2) The number of training epochs; and (3) The inference time. To provide a consistent measure of inference time across RFMP, SRFMP, and DP, we report it in terms of the number of function evaluations (NFE), which is proportional to the inference process time. Given that each function evaluation takes approximately the same time across all methods, inference time comparisons can be made directly based on NFE. For example, in real-world tasks, with an NFE of $2$, DP requires around $0.0075$\mathrm{s}$$, while RFMP and SRFMP take approximately $0.0108$\mathrm{s}$$ and $0.0113$\mathrm{s}$$, respectively. When NFE is increased to $5$, DP takes around $0.0175$\mathrm{s}$$, while RFMP and SRFMP require about $0.021$\mathrm{s}$$ and $0.0212$\mathrm{s}$$.
@@ V-B Push-T Tasks
We first consider two simple Push-T tasks, namely the Euclidean Push-T proposed in [2], which was adapted from the Block Pushing task [8], and the Sphere Push-T task, which we introduce shortly. The goal of the Euclidean Push-T task, illustrated in Figure 6, is to push a gray T-shaped object to the designated green target area with a blue circular agent. The agent’s movement is constrained by a light gray square boundary. Each observation ${\bm{o}}$ is composed of the $96\times 96$ RGB image of the current scene and the agent’s state information. We introduce the Sphere Push-T task, visualized in Figure 6, to evaluate the performance of our models on the sphere manifold. Its environment is obtained by projecting the Euclidean Push-T environment on one half of a $2$-dimensional sphere $\mathcal{S}^{2}$ of radius of $1$. This is achieved by projecting the environment, normalized to a range $[-1.5,1.5]$, from the plane $z=1$ to the sphere via a stereographic projection. The target area, the T-shaped object, and the agent then lie and evolve on the sphere. As in the Euclidean case, each observation ${\bm{o}}$ is composed of the $96\times 96$ RGB image of the current scene and the agent’s state information on the manifold. All models (i.e., RFMP, SRFMP, DP) are trained for $300$ epochs in both settings. During testing, we choose the best validation epoch and only roll out $500$ steps in the environment with an early stop rule terminating the execution when the coverage area is over $95\%$ of the green target area. The score for both Euclidean and Sphere Push-T tasks is the maximum coverage ratio during execution. The tests are performed with $50$ different initial states not present in the training set.
@@ V-B1 Euclidean Push-T
TABLE III: Euclidean Push-T: Impact of NFE on policies
@@
TABLE IV: RFMP Hyperparameters Ablation on Euclidean Push-T
@@
TABLE V: SRFMP Hyperparameters Ablation on Euclidean Push-T
@@
$\lambda_{{\bm{x}}}\quad\lambda_{\tau}$ $1$ $0.2$ $0.74$ $0.614$ $0.494$ $0.457$
$1$ $1$ $0.85$ $0.777$ $0.608$ $0.416$
$1$ $6$ $0.87$ $0.768$ $0.753$ $0.812$
@@ V-B2 Sphere Push-T
Next, we test the ability of RFMP and SRFMP to generate motions on non-Euclidean manifolds with the Sphere Push-T task. We evaluate two types of Riemannian prior distributions for RFMP and SRFMP, namely a spherical uniform distribution and wrapped Gaussian distribution. We additionally consider a Euclidean Gaussian distribution for DP. Notice that the actions generated by DP are normalized in a post-processing step to ensure that they belong to the sphere. The corresponding performance are reported in Table VI. Our results indicate that the choice of prior distribution significantly impacts the performance of both RFMP and SRFMP. Specifically, we observe that RFMP and SRFMP with a uniform sphere distribution consistently outperform their counterparts with wrapped Gaussian distribution. We hypothesize that RFMP or SRFMP benefit from having samples that are close to the data support, which leads to simpler vector fields to learn. In other words, uniform distribution provides more samples around the data distribution, which potentially lead to simpler vector fields. DP exhibit poor performance with sphere-based prior distributions, suggesting its ineffectiveness in handling such priors. Instead, DP’s performance drastically improves when using a Euclidean Gaussian distribution and higher NFE. Note that this high performance does not scale to higher dimensional settings as already evident in the real-world experiments reported in Section V-E, where the effect of ignoring the geometry of the parameters exacerbates, which is a known issue when naively operating with Riemannian data [66]. Importantly, SRFMP is consistently more robust to NFE and achieves high performance with a single NFE, leading to shorter inference times for similar performance compared to RFMP and DP.
TABLE VI: Impact of the prior distribution on Sphere Push-T Task
@@
TABLE VII: Influence of Integration Time on RFMP and SRFMP
@@
Figure 7: Action series generated by RFMP (first and third rows) and SRFMP (second and forth rows) trained on the Push-T tasks at integration times $t=\{0.8,1.2,1.6\}$.
@@ V-B3 Influence of Integration Time Boundary
We further assess the robustness of SRFMP to varying time boundaries on the Push-T tasks by increasing the time boundary during inference. The performance of both RFMP and SRFMP is summarized in Table VII with result presented for $\text{NFE}=3$ under the time boundaries $t=1$ and $t=1.2$, as well as for $\text{NFE}=4$ under the time boundaries $t=1.6$. Our results show that the performance of RFMP is highly sensitive to the time boundary, gradually declining as the boundary increases. In contrast, SRFMP demonstrates remarkable robustness, with minimal variation across different time boundaries. As illustrated in Figure 7, the quality of action series generated by RFMP noticeably deteriorates with increasing time boundaries, whereas SRFMP consistently delivers high-quality action series regardless of the time boundary.
@@ V-C Robomimic Benchmark
Next, we evaluate RFMP and SRFMP on the well-known Robomimic robotic manipulation benchmark [10]. This benchmark consists of five tasks with varying difficulty levels. The benchmark provides two types of demonstrations, namely proficient human (PH) high-quality teleoperated demonstrations, and mixed human (MH) demonstrations. Each demonstration contains multi-modal observations, including state information, images, and depth data. We report results on five tasks (Lift, Can, Square, Tool Hang, and Transport) from the Robomimic dataset with $200$ PH demonstrations for training for both state- and vision-based observations. Note that the difficulty of the selected tasks becomes progressively more challenging. The score of each of the $50$ trials is determined by whether the task is completed successfully after a given number of steps ($300$ for Lift, $500$ for Can and Square, $700$ for Tool Hang, and $500$ for Transport. The performance is then the percentage of successful trials.
Figure 8: Success rate (mean and standard deviation) on Robomimic tasks with state-based observations at different checkpoints. The models performance of Lift, Can, and Square tasks is checked every $10$ epochs throughout the $50$-epoch training process using $3$ NFE. For the Tool Hang task, the models are trained over $100$ epochs and checked every $20$ epochs using $10$ NFE.

### v3

@@ V Experiments
TABLE I: Hyperparameters for all experiments: Resolution of original and cropped image, number of parameters for the learned vector field (VF) using RFMP and SRFMP, number of ResNet parameters, training epochs, and batch size.
Experiment Image res. Crop res. RFMP VF # params SRFMP VF # params ResNet # params Epochs Batch size
Push-T Tasks
Euclidean Push-T $96\times 96$ $84\times 84$ $8.0\text{\times}{10}^{07}$ $8.14\text{\times}{10}^{07}$ $1.12\text{\times}{10}^{07}$ $300$ $256$
Sphere Push-T $100\times 100$ $84\times 84$ $8.0\text{\times}{10}^{07}$ $8.14\text{\times}{10}^{07}$ $1.12\text{\times}{10}^{07}$ $300$ $256$
State-based simulation tasks
Robomimic Lift N.A. N.A. $58\text{\times}{10}^{07} 6.6$$68\text{\times}{10}^{07}$ N.A. 50 256
Robomimic Can N.A. N.A. $58\text{\times}{10}^{07} 6.6$$68\text{\times}{10}^{07}$ N.A. $50$ $256$
Robomimic Square N.A. N.A. $58\text{\times}{10}^{07} 6.6$$68\text{\times}{10}^{07}$ N.A. $50$ $512$
Robomimic Tool Hang N.A. N.A. $58\text{\times}{10}^{07} 6.6$$68\text{\times}{10}^{07}$ N.A. $100$ $512$
Franka Kitchen N.A. N.A. $6.69\text{\times}{10}^{07}$ $6.89\text{\times}{10}^{07}$ N.A. 500 128
Vision-based simulation tasks
Robomimic Lift $2\times 84\times 84$ $2\times 76\times 76$ $9.48\text{\times}{10}^{07}$ $9.69\text{\times}{10}^{07}$ $2\times$1.12\text{\times}{10}^{07}$\$ 100 256
Robomimic Can $2\times 84\times 84$ $2\times 76\times 76$ $9.48\text{\times}{10}^{07}$ $9.69\text{\times}{10}^{07}$ $2\times$1.12\text{\times}{10}^{07}$\$ 100 256
Robomimic Square $2\times 84\times 84$ $2\times 76\times 76$ $9.48\text{\times}{10}^{07}$ $9.69\text{\times}{10}^{07}$ $2\times$1.12\text{\times}{10}^{07}$\$ 100 512
Robomimic Transport $4\times 84\times 84$ $4\times 76\times 76$ $1.24\text{\times}{10}^{08}$ $1.27\text{\times}{10}^{08}$ $4\times$1.12\text{\times}{10}^{07}$\$ 200 256
Real-world experiments
Pick & place $320\times 240$ $288\times 216$ $2.51\text{\times}{10}^{07}$ $2.66\text{\times}{10}^{07}$ $1.12\text{\times}{10}^{07}$ $300$ $256$
Mug flipping $320\times 240$ $256\times 192$ $2.51\text{\times}{10}^{07}$ $2.66\text{\times}{10}^{07}$ $1.12\text{\times}{10}^{07}$ $300$ $256$
@@ V-A1 RFMP and SRFMP Implementation
Our RFMP implementation builds on the RFM framework from Chen and Lipman [16]. We parameterize the vector field $v_{t}(\bm{a}|\bm{o};\bm{\theta})$ using the UNet architecture employed in DP [2], which consists of $3$ layers with downsampling dimensions of $(256,512,1024)$ and $(128,256,512)$ for simulated and real-world tasks. Each layer employs a $1$-dimensional convolutional residual network as proposed in [29]. We implement a Feature-wise Linear Modulation (FiLM) [59] to incorporate the observation condition vector $\bm{o}$} and time step $t$ into the UNet. Instead of directly feeding the FM time step $t$ as a conditional variable, we first project it into a higher-dimensional space using a sinusoidal embedding module, similarly to DP. For tasks with image-based observations, we leverage the same vision perception backbone as in DP [2]. Namely, we use a standard ResNet-$18$ in which we replace: (1) the global average pooling with a spatial softmax pooling, and (2) BatchNorm with GroupNorm. Our SRFMP implementation builds on the SFM framework [20]. We implement the same UNet as RFMP to represent $v_{{\bm{x}}}$ by replacing the time step $t$ by the temperature parameter $\tau$. We introduce an additional Multi-Layer Perceptron (MLP) to learn $v_{\tau}$. As for $t$ in RFMP, we employed a sinusoidal embedding for the input $\tau$. The boundaries $\tau_{0}$ and $\tau_{1}$ are set to 0 and $1$ in all experiments.
Figure 5: 2D visualization of prior distributions on the sphere: wrapped Gaussian distribution (left) and sphere uniform distribution (right). The sphere Gaussian distribution is obtained by first sampling from a Euclidean Gaussian distribution on a tangent space of the sphere (the red points), followed by the exponential map, which projects the samples onto the sphere manifold $\mathcal{S}^{d}$ (blue points). The spherical uniform distribution is computed by normalizing samples from a zero-mean Euclidean Gaussian distribution.
@@
We implement different prior distributions for different tasks. For the Euclidean Push-T and the Robomimic tasks, the action space is $\mathbb{R}^{d}$ and we thus define the prior distribution as a Euclidean Gaussian distribution for both RFMP and SRFMP. For the Sphere Push-T, the action space is the hypersphere $\mathcal{S}^{2}$. In this case, we test two types of Riemannian prior distribution, namely a spherical uniform distribution, and a wrapped Gaussian distribution [60, 54], illustrated in Figure 5. For the Franka kitchen benchmark, we define the action space as the product of manifolds $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{2}$, whose components represent the end-effector position, the end-effector orientation (encoded as quaternions), and the gripper fingers position. Regarding the real robot tasks, the action space is defined similarly as the product of manifolds $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{1}$, whose components represent the position, orientation (encoded as quaternions), and opening of the gripper. For the Franka kitchen benchmark and the real robot tasks, the Euclidean and hypersphere parts employ Euclidean Gaussian distributions and a wrapped Gaussian distribution, respectively. Notice that this choice of prior distributions follows common practice in flow matching literature, where Gaussian and uniform priors are widely adopted due to their simplicity and effectiveness [16]. While more sophisticated approaches could be explored, e.g., using Gaussian Process as prior distributions [61], or learning task-dependent priors as in D-Flow [62], the choice depends on the problem at hand. We focus here on standard priors to maintain simplicity and consistency across tasks, while isolating their impact on our framework.
@@ V-A2 Baselines
TABLE II: Prediction horizon and action space dimensionality.
@@
Euclidean Push-T $16$ 22 $32$
Sphere Push-T $16$ 22 $32$
Robomimic tasks $16$ 77 $112$
Franka Kitchen $216$ 99 $1944$
Real-world experiments $16$ 77 $112$
@@ V-A3 Evaluation methodology
We evaluate all policies using three key metrics: (1) The performance, computed as the average task-depending score across all trials, with $50$ trials for each simulated task, and $10$ trials for each real-world task; (2) The number of training epochs; and (3) The inference time. To provide a consistent measure of inference time across RFMP, SRFMP, and DP, we report it in terms of the number of function evaluations (NFE), which is proportional to the inference process time. Given that each function evaluation takes approximately the same time across all methods, inference time comparisons can be made directly based on NFE. For example, in real-world tasks, with an NFE of $2$, DP requires around $0.0075$\mathrm{s}$\$, while RFMP and SRFMP take approximately $0.0108$\mathrm{s}$\$ and $0.0113$\mathrm{s}$\$, respectively. When NFE is increased to $5$, DP takes around $0.0175$\mathrm{s}$\$, while RFMP and SRFMP require about $0.021$\mathrm{s}$\$ and $0.0212$\mathrm{s}$\$.
@@ V-B Push-T Tasks
We first consider two simple Push-T tasks, namely the Euclidean Push-T proposed in [2], which was adapted from the Block Pushing task [8], and the Sphere Push-T task, which we introduce shortly. The goal of the Euclidean Push-T task, illustrated in Figure 6, is to push a gray T-shaped object to the designated green target area with a blue circular agent. The agent’s movement is constrained by a light gray square boundary. Each observation $\bm{o}$} is composed of the $96\times 96$ RGB image of the current scene and the agent’s state information. We introduce the Sphere Push-T task, visualized in Figure 6, to evaluate the performance of our models on the sphere manifold. Its environment is obtained by projecting the Euclidean Push-T environment on one half of a $2$-dimensional sphere $\mathcal{S}^{2}$ of radius of $1$. This is achieved by projecting the environment, normalized to a range [1.5,1.5][-1.5,1.5], from the plane z=1z\!=\!1 to the sphere via a stereographic projection. The target area, the T-shaped object, and the agent then lie and evolve on the sphere. As in the Euclidean case, each observation $\bm{o}$} is composed of the $96\times 96$ RGB image of the current scene and the agent’s state information on the manifold. All models (i.e., RFMP, SRFMP, DP) are trained for $300$ epochs in both settings. During testing, we choose the best validation epoch and roll out $500$ steps in the environment with an early stop rule terminating the execution when the coverage area is over $95\%$ of the green target area. The score for both Euclidean and Sphere Push-T tasks is the maximum coverage ratio during execution. The tests are performed with $50$ different initial states not present in the training set.
@@ V-B1 Euclidean Push-T
TABLE III: Euclidean Push-T: Impact of NFE on policies.
@@
TABLE IV: RFMP Hyperparameters Ablation on Euclidean Push-T.
@@
TABLE V: SRFMP Hyperparameters Ablation on Euclidean Push-T.
@@
$\lambda_{{\bm{x}}}\quad\lambda_{\tau}$ 11 $0.2$ $0.74$ $0.614$ $0.494$ $0.457$
11 11 $0.85$ $0.777$ $0.608$ $0.416$
11 66 $0.87$ $0.768$ $0.753$ $0.812$
@@ V-B2 Sphere Push-T
Next, we test the ability of RFMP and SRFMP to generate motions on non-Euclidean manifolds with the Sphere Push-T task. We evaluate two types of Riemannian prior distributions for RFMP and SRFMP, namely a spherical uniform distribution and wrapped Gaussian distribution (see Figure 5). We additionally consider a Euclidean Gaussian distribution for DP. Notice that the actions generated by DP are normalized in a post-processing step to ensure that they belong to the sphere. The corresponding performance are reported in Table VI. Our results indicate that the choice of prior distribution significantly impacts the performance of both RFMP and SRFMP. Specifically, we observe that RFMP and SRFMP with a uniform sphere distribution consistently outperform their counterparts with wrapped Gaussian distribution. We hypothesize that RFMP or SRFMP benefit from having samples that are close to the data support, which leads to simpler vector fields to learn. In other words, uniform distribution provides more samples around the data distribution, which potentially lead to simpler vector fields. DP exhibit poor performance with sphere-based prior distributions, suggesting its ineffectiveness in handling such priors. Instead, DP’s performance drastically improves when using a Euclidean Gaussian distribution and higher NFE. Note that this high performance does not scale to higher dimensional settings as already evident in the real-world experiments reported in Section V-E, where the effect of ignoring the geometry of the parameters exacerbates, which is a known issue when naively operating with Riemannian data [65]. Importantly, SRFMP is consistently more robust to NFE and achieves high performance with a single NFE, leading to shorter inference times for similar performance compared to RFMP and DP.
TABLE VI: Impact of the prior distribution on Sphere Push-T Task.
@@
TABLE VII: Influence of Integration Time on RFMP and SRFMP.
@@
Figure 7: Action series generated by RFMP (first and third rows) and SRFMP (second and fourth rows) trained on the Push-T tasks at integration times $t=\{0.8,1.2,1.6\}$.
@@ V-B3 Influence of Integration Time Boundary
We further assess the robustness of SRFMP to varying time boundaries on the Push-T tasks by increasing the time boundary during inference. The performance of both RFMP and SRFMP is summarized in Table VII with result presented for $\text{NFE}\!=\!3$ under the time boundaries t=1t\!=\!1 and t=1.2t\!=\!1.2, as well as for $\text{NFE}=4$ under the time boundaries t=1.6t\!=\!1.6. Our results show that the performance of RFMP is highly sensitive to the time boundary, gradually declining as the boundary increases. In contrast, SRFMP demonstrates remarkable robustness, with minimal variation across different time boundaries. As illustrated in Figure 7, the quality of action series generated by RFMP noticeably deteriorates with increasing time boundaries, whereas SRFMP consistently delivers high-quality action series regardless of the time boundary.
@@ V-C Robomimic Benchmark
Next, we evaluate RFMP and SRFMP on the well-known Robomimic robotic manipulation benchmark [10]. This benchmark consists of five tasks with varying difficulty levels for which it provides two types of demonstrations, namely proficient human (PH) high-quality teleoperated demonstrations, and mixed human (MH) demonstrations. Each demonstration contains multi-modal observations, including state information, images, and depth data. We report results on five tasks (Lift, Can, Square, Tool Hang, and Transport) from the Robomimic dataset with $200$ PH demonstrations for training for both state- and vision-based observations. Note that the difficulty of the selected tasks becomes progressively more challenging. The score of each of the $50$ trials is determined by whether the task is completed successfully after a given number of steps ($300$ for Lift, $500$ for Can and Square, $700$ for Tool Hang, and $500$ for Transport). The performance is then the percentage of successful trials.
Figure 8: Success rate (mean and standard deviation) on Robomimic tasks with state-based observations at different checkpoints. The models performance of Lift, Can, and Square tasks is checked every $10$ epochs throughout the $50$-epoch training process using $3$ NFE. For the Tool Hang task, the models are trained over $100$ epochs and checked every 2020 epochs using $10$ NFE.